# OGLE-2015-BLG-0845L: A low-mass M dwarf from the microlensing parallax and xallarap effects

### v1

@@ Abstract
We present the analysis of the microlensing event OGLE-2015-BLG-0845, which was affected by both the microlensing parallax and xallarap effects. The former was detected via the simultaneous observations from the ground and Spitzer, and the latter was caused by the orbital motion of the source star in a relatively close binary. The combination of these two effects led to a direct mass measurement of the lens object, revealing a low-mass ($0.14\pm 0.05M_{\sun}$) M-dwarf at the bulge distance ($7.6\pm 1.0$ kpc). The source binary consists of a late F-type subgiant and a K-type dwarf of $\sim 1.2M_{\sun}$ and $\sim 0.9M_{\sun}$, respectively, and the orbital period is $70\pm 10$ days. OGLE-2015-BLG-0845 is the first single-lens event in which the lens mass is measured via the binarity of the source. Given the abundance of binary systems as potential microlensing sources, the xallarap effect may not be a rare phenomenon. Our work thus highlights the application of the xallarap effect in the mass determination of microlenses, and the same method can be used to identify isolated dark lenses.
@@
pubyear: 2024pagerange: OGLE-2015-BLG-0845L: A low-mass M dwarf from the microlensing parallax and xallarap effectsA
@@ 1 Introduction
The microlensing effect (Einstein, 1936; Paczynski, 1986) is able to measure the mass of faint or even dark objects (or stellar systems) independent from their flux (e.g., Gould, 1992). So far, the microlensing technique has been used to determine the masses of dozens of isolated objects, including brown dwarfs (e.g., Gould et al., 2009; Zhu et al., 2016; Shvartzvald et al., 2019), low-mass stars (e.g., Chung et al., 2017; Zhu et al., 2017b; Shin et al., 2018; Zang et al., 2020a, b), and stellar remnants (Sahu et al., 2022; Lam et al., 2022; Mroz et al., 2022), as well as another dozens of planetary systems (e.g. Gaudi et al., 2008; Bennett et al., 2015).
@@
Here $\kappa\equiv 4G/(c^{2}\rm AU)\approx 8.14{\rm mas}/M_{\odot}$ is a constant, $M_{\rm L}$ is the lens mass, and $\pi_{\rm rel}\equiv{\rm AU}(D_{\rm L}^{-1}-D_{\rm S}^{-1})$ is the lens–source relative parallax (Gould, 2000). The microlensing parallax measures the relative parallactic motion scaled to the angular Einstein radius (Gould, 1992)
@@
Here $a_{\mathrm{S}}$ is the semi-major axis of the motion of the source star (or the photocentre of the source binary) around the barycentre of the source binary system, $D_{\mathrm{S}}$ is the distance of the source from Earth, and $\hat{r}_{\mathrm{E}}$ is the projected Einstein radius in the source plane. Once $\xi_{\mathrm{E}}$ and $a_{\rm S}$ are measured, the xallarap effect provides another way to measure $\theta_{\rm E}$.
@@ 2 Observations
Even OGLE-2015-BLG-0845 was also observed by the Microlensing Observations in Astrophysics (MOA, Bond et al., 2001; Sako et al., 2008). It is called MOA-2015-BLG-277 in the MOA database. Follow-up observations from various small telescopes were also taken for the purpose of detecting planetary signals. Because of the unknown systematics and/or the relatively short time baseline, the MOA data and the follow-up data are not included in the modeling of the parallax and xallarap signals, but they provide useful constraints on the binary lens modeling (see Appendix A).
@@ 3 Modeling
The standard Paczynski (1986) curve cannot fit the light curve of OGLE-2015-BLG-0845 very well, as shown in Figure 2. While these deviations could be modeled by the annual parallax effect reasonably well, it was soon recognized that the parallax parameters from the annual parallax effect did not match those from the satellite parallax effect (Section 3.1). This led to the inclusion of additional effects, including the xallarap effect (Section 3.2). The binary lens, as well as the static binary source model, was also considered, although it could not fit the data as well as the xallarap model (see Appendix A).
@@ 3.1 The Tension Between 1L1S Annual and Space-based Parallax Model
$\boldsymbol{\pi}_{\mathrm{E}}=\frac{\pi_{\mathrm{rel}}}{\theta_{\mathrm{E}}}\frac{\boldsymbol{\mu}_{\mathrm{rel}}}{\mu_{\mathrm{rel}}},\quad\pi_{\mathrm{rel}}=\mathrm{AU}\left(\frac{1}{D_{\mathrm{L}}}-\frac{1}{D_{\mathrm{S}}}\right).$ (9)
Here $\pi_{\rm{rel}}$ and $\vec{\mu}_{\rm{rel}}$ are the relative parallax and the relative proper motion between the lens and source, respectively (Gould, 2000).
@@
The xallarap effect is the reflection of the orbital motion of the source on the light curve, as is shown in Figure 1. In principle, one should also consider the companion star (hereafter the secondary source) in the light curve modeling. However, the probability is high that the secondary source only contributes a marginal fraction of the total source flux because of the steep relation between stellar mass and luminosity. Therefore, we begin the xallarap modeling by including only one luminous source. For simplicity, the orbit of the source star is assumed to be circular, and the impact of an elliptical orbit is discussed in the section 5.1.
@@
Here $i_{\xi}$ is the orbital inclination, $\phi_{\xi}$ is the phase of the source relative to the ascending node at the reference epoch $t_{\rm ref}$, and $\theta_{\xi}$ measures the ascending node relative to the lens–source relative trajectory projected onto the source plane. The amplitude of the xallarap motion, $\xi_{\rm E}$, the semi-major axis of the binary motion scaled to the projected Einstein ring radius, as is defined in Equation 3.
@@
The binary motion of the source, $\boldsymbol{S}(t)$, is given by the circular motion starting from a zero phase with an amplitude of unity, and the velocity vector, $\boldsymbol{v_{\rm ref}}$, is the instantaneous velocity of the source at reference epoch $t_{\rm ref}$.
@@ 4.1 Source Properties
$\frac{a_{\rm S}}{\rm AU}=\frac{q}{(1+q)^{2/3}}\left(\frac{M_{\rm S}}{M_{\odot}}\right)^{1/3}\left(\frac{P_{\xi}}{\rm yr}\right)^{2/3}$ (19)
Here $M_{\rm S}$ is the masses of the primary and $q_{\xi}$ is the mass ratio. The values of these masses are key in determining $\theta_{\rm E}$, as the xallarap amplitude and the source binary period are both measured in the xallarap modeling.
@@
$(V-I,I)_{\rm{S2}}=(1.4\pm 0.3,20.6\pm 0.5)$ (21)
for the secondary source. The above values are derived from the (+,)(+,-)( + , - ) solution, i.e., the minimum $\chi^{2}$ solution, but the difference among the four degenerate solutions in these numbers is statistically negligible.
@@
following the method of Nataf et al. (2013). At the location of OGLE-2015-BLG-0845 , the intrinsic colour and dereddened magnitude of the red clump are
@@
The exact values and uncertainties of these parameters are included in Table 2 for all four degenerate solutions.
@@ 4.2 Mass and Distance of the Lens
Here $\tau_{\rm{dust}}=0.1\rm{kpc}/\sin(b)$ is the scale length of the dust towards the galactic bulge, with $b$ the Galactic latitude of the source system. For all four solutions, the lens mass and distance, derived from the parallax and angular Einstein radius measurements, are well within the allowed region by the lens flux upper limit.
@@ 5.2 Source Binary Confirmation
With a baseline magnitude of $I\approx 18.5$ and almost zero blending, the source binary of OGLE-2015-BLG-0845 is potentially accessible by spectroscopic observations from the ground. Such observations can not only confirm but also further refine the derived properties of the source binary (and thus the lens object).
We observed the source with the MagE spectrograph (Marshall et al., 2008) on the 6.5 m Magellan Baade Telescope at Las Campanas Observatory on July 2, 2023. Three exposures, each with 20 min, were taken, and the resulting spectrum has a signal-to-noise ratio (S/N) per resolution of about five. Given the flux ratio of $\sim 0.15$, we expect to see only a single component in the spectra. We compared the observed spectrum with synthetic spectra generated with a stellar atmosphere grid MARCS (Gustafsson et al., 2008) and a radiative transfer code turbospectrum (Plez, 2012) integrated together by iSpec (Blanco-Cuaresma, 2019). The observed spectrum can be fitted with the spectra of late F-type subgiants and dwarfs reasonably well, while it disfavors the early F-type as well as the typical G-type stars. This is in agreement with the CMD fitting result.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (grant No. 12133005 & 12173021) and the CASSACA grant CCJRF2105. We also acknowledge the Tsinghua Astrophysics High-Performance Computing platform for providing computational and data storage resources.
@@ Appendix A Binary Lens Interpretation
An additional companion to the lens object may also lead to distortions in the microlensing light curve (Mao & Paczynski, 1991; Gould & Loeb, 1992). As long as the source star stays relatively away from the caustics of the binary lens, such distortions are not prominent features and may resemble those produced by the xallarap effect (e.g., Rota et al., 2021; Yang et al., 2024). We therefore carry out 2L1S modeling to make sure the detected signal is indeed caused by the xallarap effect.
In this 2L1S modeling, we exclude the Spitzer data but include the MOA data. The latter has a higher cadence, which can be useful in constraining the short-timescale signals arising from the binary lens. To describe the 2L1S model, four new parameters are introduced, namely $\rho$, $\alpha$, $s$, and $q$. Here, $\rho$ represents the angular size of the source normalized to the angular Einstein radius, $\alpha$ is the angle between the source trajectory and the binary axis, and $s$ and $q$ are the projected separation and mass ratio between the two lens components, respectively. Following a thorough grid search and refined MCMC sampling of the posterior distribution, we only identify one binary lens solution that matches the observed data the best. This best-fit 2L1S model is shown in Figure 10, and the parameter values are given in Table 3 for the purpose of completeness. As shown in Figure 10, the best-fit 2L1S model is not able to explain fully the deviations in the ground-based data, and it is worse than the 1L2S xallarap model by $\Delta\chi^{2}$ of $\sim 130$. We therefore exclude the binary lens model as a plausible solution for OGLE-2015-BLG-0845 .
Table 3: Parameters of 1L2S xallarap + parallax solutions for OGLE-2015-BLG-0845
@@
$t_{0}$ (HJD) $7199.3746\pm 0.0065$
$u_{0}$ $0.06632\pm 0.00079$
$t_{\mathrm{E}}$ (days) $31.80\pm 0.32$
$\rho$ $7207.56\pm 0.52$
$\alpha$ $0.224\pm 0.015$
$s$ $0.136\pm 0.013$
$q$ $0.139\pm 0.013$
Blend $-0.259\pm 0.023$
$\chi^{2}$/dof 2761.35/2836
@@
Figure 10: The comparison between the 2L1S model and the 1L2S xallarap + parallax model. The figure description is similar to that of Figure 4. In the figure, the MOA data is binned with a one-day bin size, while the fitting process does not involve any binning. The inset in the upper panel shows the trajectory of the source as well as the caustic curve of the binary lens and the position of the lens.

### v2

@@ Abstract
We present the analysis of the microlensing event OGLE-2015-BLG-0845, which was affected by both the microlensing parallax and xallarap effects. The former was detected via the simultaneous observations from the ground and Spitzer, and the latter was caused by the orbital motion of the source star in a relatively close binary. The combination of these two effects led to a mass measurement of the lens object, revealing a low-mass ($0.14\pm 0.05M_{\sun}$) M-dwarf at the bulge distance ($7.6\pm 1.0$ kpc). The source binary consists of a late F-type subgiant and a K-type dwarf of $\sim 1.2M_{\sun}$ and $\sim 0.9M_{\sun}$, respectively, and the orbital period is $70\pm 10$ days. OGLE-2015-BLG-0845 is the first single-lens event in which the lens mass is measured via the binarity of the source. Given the abundance of binary systems as potential microlensing sources, the xallarap effect may not be a rare phenomenon. Our work thus highlights the application of the xallarap effect in the mass determination of microlenses, and the same method can be used to identify isolated dark lenses.
@@
pubyear: 2024pagerange: OGLE-2015-BLG-0845L: A low-mass M dwarf from the microlensing parallax and xallarap effectsB
@@ 1 Introduction
The microlensing effect (Einstein, 1936; Paczynski, 1986) is able to measure the mass of faint or even dark objects (or stellar systems) independent from their flux (e.g., Gould, 1992). So far, the microlensing technique has been used to determine the masses of dozens of isolated objects, including brown dwarfs (e.g., Gould et al., 2009; Zhu et al., 2016; Shvartzvald et al., 2019), low-mass stars (e.g., Chung et al., 2017; Zhu et al., 2017b; Shin et al., 2018; Zang et al., 2020a, b), and stellar remnants (Sahu et al., 2022; Lam et al., 2022; Mróz et al., 2022), as well as another dozens of planetary systems (e.g. Gaudi et al., 2008; Bennett et al., 2015).
@@
Here $\kappa\equiv 4G/(c^{2}\rm AU)\approx 8.14\ {\rm mas}/M_{\odot}$ is a constant, $M_{\rm L}$ is the lens mass, and $\pi_{\rm rel}\equiv{\rm AU}(D_{\rm L}^{-1}-D_{\rm S}^{-1})$ is the lens–source relative parallax (Gould, 2000). The distances to the source and to the lens are $D_{\mathrm{S}}$ and $D_{\rm L}$, respectively. The microlensing parallax measures the relative parallactic motion scaled to the angular Einstein radius (Gould, 1992)
@@
Here $a_{\mathrm{S}}$ is the semi-major axis of the motion of the source star (or the photocentre of the source binary) around the barycentre of the source binary system, and $\hat{r}_{\mathrm{E}}$ is the projected Einstein radius in the source plane. Once $\xi_{\mathrm{E}}$ and $a_{\rm S}$ are measured, the xallarap effect provides another way to measure $\theta_{\rm E}$.
@@ 2 Observations
Event OGLE-2015-BLG-0845 was also observed by the Microlensing Observations in Astrophysics (MOA, Bond et al., 2001; Sako et al., 2008). It is called MOA-2015-BLG-277 in the MOA database. Follow-up observations from various small telescopes were also taken for the purpose of detecting planetary signals. Because of the unknown systematics and/or the relatively short time baseline, the MOA data and the follow-up data are not included in the modeling of the subtle signals such as the parallax and xallarap signals, as is commonly done in the modeling of other microlensing events (e.g., Mróz et al., 2022). In the case of OGLE-2015-BLG-0845 , because there are about twice as many observations from MOA as from OGLE, the more informative OGLE data would be substantially downweighted if one were to include the MOA data in the modeling. We have nevertheless included the MOA data in the single source binary lens (2L1S) and binary source without xallarap effect (1L2L static) modelings, given that the potential signals from such models are relatively short-timescale (see Appendix A and Appendix B).
@@ 3 Modeling
The standard Paczynski (1986) curve cannot fit the light curve of OGLE-2015-BLG-0845 very well, as shown in Figure 2. While these deviations could be modeled by the annual parallax effect reasonably well, it was soon recognized that the parallax parameters from the annual parallax effect did not match those from the satellite parallax effect (Section 3.1). This led to the inclusion of additional effects, including the xallarap effect (Section 3.2). The binary lens, as well as the static binary source model, was also considered, although it could not fit the data as well as the xallarap model (see Appendix A and Appendix B).
@@
The color constraint, which has been frequently used to infer the source flux in Spitzer band based on the known source flux values in the ground-based bandpasses (e.g., Calchi Novati et al., 2015b; Yee et al., 2015), is not used in the present event. The color–color relation is usually constructed based on field stars of similar colors, so its applicability to binary stars is questionable. One may in principle derive two separate color–color relations for the two components of the source binary, but this is difficult to implement as the source binary is not well constrained by the ground-based data alone (see Section 3.3). Instead of such statistical test as the color constraint, we will use more direct tests to verify the validity of the Spitzer data.
@@ 3.1 The Tension Between 1L1S Annual and Space-based Parallax Model
$\boldsymbol{\pi}_{\mathrm{E}}=\frac{\pi_{\mathrm{rel}}}{\theta_{\mathrm{E}}}\frac{\boldsymbol{\mu}_{\mathrm{rel}}}{\mu_{\mathrm{rel}}}.$ (9)
Here $\vec{\mu}_{\rm{rel}}$ is the relative proper motion between the lens and source, respectively (Gould, 2000).
@@
The xallarap effect is the reflection of the orbital motion of the source star/system on the light curve, as is shown in Figure 1. Whenever the xallarap effect is invoked, it almost always assumes that the microlensing source is a binary (or higher multiple). By including only one bright source in the xallarap modeling, we have assumed that the source binary is unresolved or that the secondary component is very faint/dark, which will be checked at the end of this subsection.  [^2] In microlensing, a binary source being unresolved means that $\xi_{\rm E}\ll\theta_{\rm E}$ and $\xi_{\rm E}\ll u_{0}$. In either case, the “source” here refers to the photocenter of the binary system rather than any single star. For simplicity, the orbit of the source is assumed to be circular, and the impact of an elliptical orbit is discussed in the section 5.1.
@@
Here $i_{\xi}$ is the orbital inclination, $\phi_{\xi}$ is the phase of the source relative to the ascending node at the reference epoch $t_{\rm ref}$, and $\theta_{\xi}$ measures the ascending node relative to the lens–source relative trajectory projected onto the source plane. The amplitude of the xallarap motion, $\xi_{\rm E}$, is the semi-major axis of the binary motion scaled to the projected Einstein ring radius, as is defined in Equation 3.
@@
The binary motion of the source, $\boldsymbol{S}(t)$, is given by the circular motion starting from a zero phase with an amplitude of unity, and the velocity vector, $\boldsymbol{v_{\rm ref}}$, is the instantaneous velocity of the source relative to the center of mass of the source binary at reference epoch $t_{\rm ref}$.
@@
The above xallarap solutions are identified through a free search in the whole parameter space. To ensure that we have not missed any degenerate solutions, a grid search in the orbital period of the source binary has been performed. In total 39 period values evenly spaced in logarithmic scale from 7 to 700 days are selected, for each chosen period, we fix $P_{\xi}$ to the grid value and the parallax parameters to the globally best-fit values and search for the optimal parameter set that minimizes the model $\chi^{2}$ via MCMC. The results are shown in Figure 6 for all four solutions. Within the period range that was searched, we find no other solution that can fit the joint OGLE+Spitzer data equally well ($\Delta\chi^{2}\leq 100$). Therefore, the solutions presented in Table 1 are the only viable xallarap solutions of OGLE-2015-BLG-0845 .
Figure 6: The change of $\Delta\chi^{2}$ versus $P_{\xi}$ of the four degenerate solutions. The $\chi^{2}_{\rm min}$ is the best-fitting $\chi^{2}$ of each solution, and the corresponding period is shown as the red solid line. The gray dashed line indicates the 1-year period.
@@ 4.1 Source Properties
$\frac{a_{\rm S}}{\rm AU}=\frac{q_{\xi}}{(1+q_{\xi})^{2/3}}\left(\frac{M_{\rm S1}}{M_{\odot}}\right)^{1/3}\left(\frac{P_{\xi}}{\rm yr}\right)^{2/3}$ (19)
Here $M_{\rm S1}$ is the mass of the primary. The values of these masses are key in determining $\theta_{\rm E}$, as the xallarap amplitude and the source binary period are both measured in the xallarap modeling.
@@
$(V-I,I)_{\rm{S2}}=(1.5\pm 0.4,20.9\pm 0.5)$ (21)
for the secondary source. The above values are derived from the (,)(-,-)( - , - ) solution, i.e., the minimum $\chi^{2}$ solution, but the difference among the four degenerate solutions in these numbers is statistically negligible.
Figure 7: CMD for the $2\arcmin\times 2\arcmin$ square centred on OGLE-2015-BLG-0845 . The best-fitting centroid of the red clump is shown as the red dot. The position of the S1 and S2 from solution (,)(-,-)( - , - ) are also marked on the figure. The blending is not shown, because it is not well constrained and generally consistent with zero. The relative positions of other degenerate solutions are the same as shown in Figure 9.
@@
following the method of Nataf et al. (2013). We show the position of the red clumps centre as well as the S1, S2 from the (,)(-,-)( - , - ) solution in Figure 7. At the location of OGLE-2015-BLG-0845 , the intrinsic colour and dereddened magnitude of the red clump are
@@
The exact values and uncertainties of these parameters are included in Table 2 for all four degenerate solutions. For OGLE-2015-BLG-0845 we use $D_{\rm S}=8.2\pm 1.4$ kpc, which is the distance and the radial extension of the bulge at the given galactic direction (Nataf et al., 2013).
@@ 4.2 Mass and Distance of the Lens
Here $\tau_{\rm{dust}}=0.1{\rm kpc}/\sin(|b|)$ is the scale length of the dust towards the galactic bulge, with $b$ the Galactic latitude of the source system. For all four solutions, the lens mass and distance, derived from the parallax and angular Einstein radius measurements, are well within the allowed region by the lens flux upper limit.
@@ 5.2 Source Binary Confirmation
With a baseline magnitude of $I\approx 18.5$ and a blending fraction that is low ($\lesssim 0.1$) and consistent with zero (Table 1), the source binary of OGLE-2015-BLG-0845 is potentially accessible by spectroscopic observations from the ground. Such observations can not only confirm but also further refine the derived properties of the source binary (and thus the lens object).
[…]
Figure 11: The observed MagE spectrum as well as the best-fitting stellar model. The MagE observation and best-fitting model are shown in blue and orange solid lines, respectively.
@@ Acknowledgements
We thank the reviewer for useful comments and suggestions. This work is supported by the National Natural Science Foundation of China (grant No. 12133005 & 12173021) and the CASSACA grant CCJRF2105. R. P. has been supported by the National Science Center, Poland, grant No. 2021/42/E/ST9/00038. For the purpose of Open Access, the author has applied a CC-BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission. We also acknowledge the Tsinghua Astrophysics High-Performance Computing platform for providing computational and data storage resources.
@@ Appendix A Binary Lens Interpretation
An additional companion to the lens object may also lead to distortions in the microlensing light curve (Mao & Paczynski, 1991; Gould & Loeb, 1992). As long as the source star stays relatively away from the caustics of the binary lens, such distortions are not prominent features and may resemble those produced by the xallarap effect (e.g., Rota et al., 2021; Yang et al., 2024). We therefore carry out 2L1S modeling including the lens orbital motion to make sure the detected signal is indeed caused by the xallarap effect.
In this 2L1S modeling, we have included the MOA data as well based on the arguments in Section 2. To describe the 2L1S model, six new parameters are introduced in addition to the 1L1S parameters, including four standard parameters ($\rho$, $\alpha$, $s$, and $q$) and two that describe the linearized lens orbital motion (${\rm d}\alpha/{\rm d}t$ and ${\rm d}s/{\rm d}t$). Here, $\rho$ represents the angular size of the source normalized to the angular Einstein radius, $\alpha$ is the angle between the source trajectory and the binary axis, and $s$ and $q$ are the projected separation and mass ratio between the two lens components, respectively. The orbital motion of the lens system is included in order to have a fair comparison with the binary source model with the xallarap effect. Following a thorough grid search and refined MCMC sampling of the posterior distribution, we only identify one binary lens solution that matches the observed data the best. This best-fit 2L1S model is shown in Figure 13, and the parameter values are given in Table 3 for the purpose of completeness. As shown in Figure 13, the best-fit 2L1S model is not able to explain the deviations in the ground-based data fully, and it is worse than the 1L2S xallarap model by $\Delta\chi^{2}$ of $\sim 230$. Here the 1L2S xallarap model is the same as given in Table 1, and we have introduced a flux ratio in the MOA band to fit the MOA data. Given the relatively large $\Delta\chi^{2}$ value, we therefore exclude the binary lens model as a plausible solution for OGLE-2015-BLG-0845 .
Table 3: Parameters of the best-fit 2L1S solution for OGLE-2015-BLG-0845 .
@@
$t_{0}$ (HJD) $7199.6348\pm 0.0087$
$u_{0}$ $0.0547\pm 0.0010$
$t_{\mathrm{E}}$ (days) $40.5\pm 0.5$
$\pi_{\mathrm{E,N}}$ $-0.0030\pm 0.0003$
$\pi_{\mathrm{E,E}}$ $0.0797\pm 0.0011$
$\rho$ $0.008\pm 0.003$
$\alpha$ ($\deg$) $7\pm 52$
$s$ $0.641\pm 0.010$
$q$ $0.0058\pm 0.0003$
${\rm d}\alpha/{\rm d}t$ (deg / yr) $-88\pm 8$
${\rm d}s/{\rm d}t$ (yr-1) $0.44\pm 0.08$
Blend Fractioni $0.089\pm 0.016$
$\chi^{2}$/dof 3095.11/2923
@@
Figure 13: The comparison between the 2L1S model and the 1L2S xallarap + parallax model. The figure description is similar to that of Figure 4. In the figure, we present both the original MOA data without error bars and the MOA data binned into daily cadence, which are denoted as green dots and circles with error bars, respectively.. The inset in the upper panel shows the trajectory of the source as well as the caustic curve of the binary lens and the position of the lens.
Appendix B Static Binary Source Interpretation
In Table 4, we present the best-fit binary source solution without orbital motion (i.e., 1L2S static) here for completeness. In the 1L2S static model, the position of the secondary source is given by $t_{0,2}$ and $u_{0,2}$, which are the time of closest approach and the impact parameter of the secondary source, respectively (e.g., Hwang et al., 2013). Similar to the 1L2S xallarap model, the flux ratios between the two sources in individual bandpasses are also required.
The comparison between the best-fit 1L2S static solution and the best 1L2S xallarap solution is shown in Figure 14, with the latter one the same as in Appendix A. Similar to the origin of the four-fold parallax degeneracy, there exist multiple degenerate solutions for the 1L2S xallarap explanation, and here we select the one with the smallest model $\chi^{2}$. As shown in Figure 14, even this best-fit 1L2S static solution is worse than the best 1L2S xallarap solution by $\Delta\chi^{2}>230$, and it fails to explain the general behaviour of the light curve between $71957160<{\rm HJD}\arcmin<$ 7195. Therefore, the 1L2S static model is excluded.
Table 4: Parameters of best-fit 1L2S static solution for OGLE-2015-BLG-0845 .
Parameters (+,)(+,-)( + , - )
$t_{0,1}$ (HJD) $7199.376\pm 0.006$
$u_{0,1}$ $0.0666\pm 0.0007$
$t_{\mathrm{E}}$ (days) $31.8\pm 0.3$
$\pi_{\mathrm{E,N}}$ $-0.0968\pm 0.0010$
$\pi_{\mathrm{E,E}}$ $0.1129\pm 0.0010$
$t_{0,2}$ (HJD) $7207.6\pm 0.4$
$u_{0,2}$ $0.230\pm 0.013$
$q_{\mathrm{f,OGLE}}$ $0.137\pm 0.011$
$q_{\mathrm{f,OGLE,V}}$ $0.106\pm 0.023$
$q_{\mathrm{f,MOA}}$ $0.137\pm 0.011$
$q_{\mathrm{f,Spitzer}}$ $0.037\pm 0.007$
Blend Fraction $-0.265\pm 0.017$
$\chi^{2}$/dof 3104.71/2921
  • NOTE. HJD=HJD-2450000.
Figure 14: The comparison between the 1L2S static and the 1L2S xallarap model. The figure description is similar to that of Figure 13.